# Conformalized Interval Arithmetic with Symmetric Calibration

### v1

@@ Abstract.
Rui Luo***ruiluo@cityu.edu.hk1 and Zhixin Zhouzhixzhou@cityu.edu.hk1
1City University of Hong Kong
1. Introduction
Conformal prediction [35] has gained popularity as a method for uncertainty quantification due to its validity guarantee. It plays a crucial role in modern decision-making processes, particularly in domains where understanding the joint distribution of random variables is essential. Under the exchangeability assumptions on the calibration and test samples, conformal prediction can provide prediction sets with valid coverage guarantees. Existing works have studied prediction intervals for single targets [16, 29, 31]. Given a black box algorithm to estimate the target label, current conformalized algorithms can provide prediction intervals for the target with reliable coverage probability. However, although numerous studies have developed the methodology of conformal prediction, they are limited to predicting a single label, leaving a gap in finding the prediction set for the sum of labels.
This gap is particularly relevant in applications such as transductive conformal prediction on traffic networks. For example, existing Graph Neural Network (GNN) methods can predict the label of each road, where the label can be considered as the cost of traversing that road. This problem has been studied in [11, 21, 19]. Conformalized GNN can output a prediction set of each edge’s label, but the cost of a route, which is the sum of the labels of the edges on the route, cannot be directly obtained by applying conformal prediction to individual edges. This challenge arises from two main aspects. First, the sum or average of random variables involves the convolution of the density function, and simple interval arithmetic, such as adding up the lower and upper bounds, cannot provide a confidence interval with the desired coverage. Second, the coverage of each confidence interval is in a marginal sense, so the coverage of two labels from two confidence intervals are dependent events. Consequently, it is difficult to use the conformal prediction set for a single label to devise a prediction set for multiple labels. To address this critical gap in the current literature on conformal prediction and expand its applicability to a broader range of uncertainty quantification problems, we introduce the method Conformal Interval Arithmetic (CIA), specifically designed to estimate the average or other symmetric functions of unknown labels over a certain index set, demonstrating the usefulness of our problem setting in many applications where people are interested in obtaining estimates about multiple labels.
@@ 1. Introduction
In addition to the methodology of symmetric calibration and the validity guarantee discussed above, we will introduce some important extensions of our method to make it more widely applicable and efficient. First, symmetric calibration is easy to understand when the groups of indices are disjoint. We relax this condition to allow for small amounts of overlap between groups. Second, given prior knowledge of the number of unknown labels being summed, we can stratify the prediction into different levels. Finally, we can incorporate other conformal prediction methods, such as conformalized quantile regression [29], to improve prediction efficiency.
@@
  3. (3)
    We apply our methods on datasets in [30], and show the reliability of our method. The code of our method will be published as open source.
2. Preliminary and Problem Setup
@@
2.1. Split Conformal Prediction
@@
We will first focus on this choice and then extend it to other choices in later sections. For data in $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{cal}}}$, we evaluate $s_{i}=|y_{i}-\widehat{y}_{i}|$ and let $Q_{1-\alpha}$ be the $\lceil(1+|\mathcal{I}_{\textrm{cal}}|)(1-\alpha)\rceil$-th smallest value of $|y_{i}-\widehat{y}_{i}|$ for $i\in\mathcal{I}_{\textrm{cal}}$. Then, for $i\in\mathcal{I}_{\textrm{test}}$, the prediction set of $y_{i}$ is defined as
@@
2.2. Problem Setup and a Naive Approach
@@
$\mathbb{P}(Y_{i}+Y_{j}\in[L_{i}+L_{j},U_{i}+U_{j}])\geq\mathbb{P}(Y_{i}\in[L_{i},U_{i}]\text{ and }Y_{j}\in[L_{j},U_{j}])\geq 1-2\alpha.$
@@
3. Conformal Interval Arithmetic
@@
$s_{k}:=s(z_{k}):=\left|\sum_{i\in S_{k}}y_{i}-\widehat{y}_{i}\right|,k\in[K]$
@@
(1) $\left[\sum_{i\in S_{K+1}}\widehat{y}_{i}-Q_{1-\alpha},\sum_{i\in S_{K+1}}\widehat{y}_{i}+Q_{1-\alpha}\right].$
@@
(2) $\mathbb{P}(Z_{1},\dots,Z_{K+1})=\mathbb{P}(Z_{\pi(x)},\dots,Z_{\pi(K+1)}),$
@@ 4. Symmetric Calibration
(3) $S^{\textrm{cal}}_{k}:=S_{k}\cap\mathcal{I}_{\textrm{cal}}\quad\text{ and }\quad S^{\textrm{test}}_{k}=S_{k}\cap\mathcal{I}_{\textrm{test}}.$
Our task becomes finding the prediction set for $\sum_{i\in S^{\textrm{test}}_{k}}y_{i}$ for every $k\in[K+1]$. However, since the algorithm is equivalent, we can focus on finding the confidence interval for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$. Using conformal interval arithmetic, we define the score function similar to (3):
(4) $s^{\textrm{cal}}_{k}=\left|\sum_{i\in S^{\textrm{cal}}_{k}}y_{i}-\widehat{y}_{i}\right|$
Let $Q_{1-\alpha}$ be the $\lceil(1+K)(1-\alpha)\rceil$-th smallest value of $s^{\textrm{cal}}_{k},k=1,\dots,K$. Then we define the prediction set for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$, denoted by $\mathcal{C}(S^{\textrm{test}}_{K+1})$:
(5) $\left[\sum_{i\in S^{\textrm{test}}_{K+1}}\widehat{y}_{i}-Q_{1-\alpha},\sum_{i\in S^{\textrm{test}}_{K+1}}\widehat{y}_{i}+Q_{1-\alpha}\right].$
@@
1:Input: labeled samples $\subscripttest\{(x_{i},y_{i})\}_{i\in\mathcal{I}\backslash\mathcal{I}_{\textrm{test}}}$, unlabeled samples $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$, a black box algorithm $\mathcal{B}$, disjoint index sets $S_{1},\dots,S_{K},S_{K+1}\subseteq\mathcal{I}$.
2:Output: Prediction sets $\mathcal{C}(S^{\textrm{test}}_{k})$ for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$.
@@
[…]
@@
6:$s^{\textrm{cal}}_{k}=\left|\sum_{i\in S^{\textrm{cal}}_{k}}y_{i}-\widehat{f}(x_{i})\right|$ for $k\in[K]$
7:$Q_{1-\alpha}\leftarrow\lceil(1+K)(1-\alpha)\rceil$-th smallest value of $\{s^{\textrm{cal}}_{k}\}_{k\in[K]}$.
@@
We will analyze the coverage probability of $\mathcal{C}(S^{\textrm{test}}_{K+1})$. In this section, we assume that $S_{1},\dots,S_{K},S_{K+1}$ are disjoint. We will introduce the reasoning behind this in two steps. Firstly, we assume that the score functions $s^{\textrm{cal}}_{k},k\in[K+1]$, defined in (3) are exchangeable. $s^{\textrm{cal}}_{K+1}$ has a uniform rank over $[K+1]$, so $s^{\textrm{cal}}_{K+1}\leq Q_{1-\alpha}$ with probability at least $1-\alpha$. Equivalently, this suggests that the interval
(6) $\left[\sum_{i\in S^{\textrm{cal}}_{K+1}}\widehat{y}_{i}-Q_{1-\alpha},\sum_{i\in S^{\textrm{cal}}_{K+1}}\widehat{y}_{i}+Q_{1-\alpha}\right]$
@@
(7) $\mathbb{P}(i\in\mathcal{I}_{\textrm{cal}})=\mathbb{P}(i\in\mathcal{I}_{\textrm{test}})=0.5$
@@
The proof will appear in the appendix in the supplimentary material. In this theorem, we have only shown the validity for $S^{\textrm{test}}_{K+1}$. However, since we assume the group exchangeability on $[K+1]$, the result can also be stated as
$\mathbb{P}\left(\sum_{i\in S^{\textrm{test}}_{k}}y_{i}\in\mathcal{C}(S^{\textrm{test}}_{k})\right)\geq 1-\alpha,$
for every $k\in[K+1]$ by changing the algorithm accordingly. To be more precise, to provide prediction interval for sum over $S^{\textrm{test}}_{k}$, $Q_{1-\alpha}$ is $\lceil(1+K)(1-\alpha)\rceil$-th smallest value of $s^{\textrm{cal}}_{\ell}$ for $\ell\in[K+1]\backslash\{k\}$. The rest of the steps remain the same as in Algorithm 1.
@@
Figure 1. Diagram illustrating the partition of data into calibration and test sets for Proposition 1 and Theorem 4.1. The sums within each of $S_{1},\dots,S_{K}$ (Top) serve as the calibration data to predict the sum within $S_{K+1}$, mirroring the calibration row for the theorem’s diagram (Bottom). As $S^{\textrm{cal}}_{K+1}$ and $S^{\textrm{test}}_{K+1}$ are exchangeable, their sums have the same distribution. Therefore, the prediction set (6) can be derived similarly to (5).
@@
The random splitting assumption (7) is slightly different from the procedure of holding out a calibration set $\mathcal{I}_{\textrm{cal}}$ with the same size as $\mathcal{I}_{\textrm{test}}$ in Algorithm 1. Theorem 4.1 suggests that it only requires the calibration set and test set have roughly the same size. Suppose $\mathcal{I}_{\textrm{cal}}$ and $\mathcal{I}_{\textrm{test}}$ are partitions of $\mathcal{I}_{\textrm{cal}}\cup\mathcal{I}_{\textrm{test}}$ with equal size, this complicates the proof of the theorem, as we can only swap $S^{\textrm{cal}}_{K+1}$ and $S^{\textrm{test}}_{K+1}$ when they have the same size.
5. Extensions
In this section, we will extend the applicability of the proposed method to handle overlapping subsets. Furthermore, we will introduce a stratified technique and utilize Conformal Quantile Regression (CQR) [29] to enhance the efficiency of our approach.
5.1. Overlapping Subsets
Algorithm 1 can be applied to overlapping subsets $S_{1},\dots,S_{K},S_{K+1}$, but the conclusion of Theorem 4.1 is valid only if the subsets are disjoint. There are applications where we must assume that the subsets overlap. We will study these applications in Section 7.2. Theoretically, overlapping subsets affect the validity of the method. The following theorem shows that as long as each subset overlaps with only a small portion of the other subsets, the effect on the coverage probability is acceptable.
Theorem 5.2.
We still assume condition (b) and (c) in Theorem 4.1, and suppose that
@@ 5.1. Overlapping Subsets
Then the coverage probability satisfies
@@
The proof will appear in the appendix in the supplimentary material. Similar as the result of Theorem 4.1, the coverage probability guarantee can be generalized to any $S^{\textrm{test}}_{k}$ for any $k\in[K+1]$.
5.3. Stratified Conformal Prediction
Given the known number of unknown labels in each test set, we can leverage this information to enhance the efficiency of conformal prediction. For instance, consider a case where $|S^{\textrm{test}}_{K+1}|$ is relatively small, such as $|S^{\textrm{test}}_{K+1}|=1$. In this case, it is more appropriate to compare the residual with other residuals that have a smaller number of unknown labels in the calibration set. By incorporating this idea, we can modify Algorithm 1 and derive a stratified version of the algorithm that takes advantage of this additional information to improve its efficiency.
@@ 5.3. Stratified Conformal Prediction
1:Input: labeled samples $\subscripttrain\{(y_{i},\widehat{y}_{i})\}_{i\in\mathcal{I}\backslash\mathcal{I}_{\textrm{train}}}$, disjoint index sets $S_{1},\dots,S_{K},S_{K+1}\subseteq\mathcal{I}$. Partition of positive integers $C_{1},C_{2},\dots$
2:Output: Prediction sets $\mathcal{C}(S^{\textrm{test}}_{k})$ for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$.
3:Find $C_{j}$ such that $|S^{\textrm{test}}_{K+1}|\in C_{j}$.
4:$s^{\textrm{cal}}_{k}=\left|\sum_{i\in S^{\textrm{cal}}_{k}}y_{i}-\widehat{y}_{i}\right|$ for $k\in[K]$.
@@
6:$Q^{j}_{1-\alpha}\leftarrow\lceil(1+n_{j})(1-\alpha)\rceil$-th smallest value of $\{s^{\textrm{cal}}_{k}:|S^{\textrm{cal}}_{k}|\in C_{j}\}$.
@@
5.4. Conformal Quantile Regression
From the theorems, we can observe that the form of the score function can be modified. For instance, we can adopt Conformal Quantile Regression (CQR) [29]. The score function can be defined as
$s^{\textrm{cal}}_{k}=\max\left\{\sum_{i\in S^{\textrm{cal}}_{k}}\widehat{y}_{i}^{\alpha/2}-y_{i},\sum_{i\in S^{\textrm{cal}}_{k}}y_{i}-\widehat{y}^{1-\alpha/2}_{i}\right\}$
for $k\in[K]$, where $\widehat{y}^{\alpha/2}_{i}$ and $\widehat{y}^{1-\alpha/2}_{i}$ are the outputs of quantile regression at quantiles $\alpha/2$ and $21-\alpha/$ 2, respectively, to predict $y_{i}$. We can obtain $Q_{1-\alpha}$ in Algorithm 1 using the score functions of CQR. The prediction set $\mathcal{C}(S^{\textrm{test}}_{K+1})$ is then given by
$\left[\sum_{i\in S^{\textrm{test}}_{K+1}}\widehat{y}^{\alpha/2}_{i}-Q_{1-\alpha},\sum_{i\in S^{\textrm{test}}_{K+1}}\widehat{y}^{1-\alpha/2}_{i}+Q_{1-\alpha}\right].$

### v2

@@ Abstract
Codehttps://github.com/luo-lorry/CIA
Extended versionhttps://arxiv.org/abs/2408.10939
1 Introduction
Conformal prediction (Vovk, Gammerman, and Shafer 2005) has gained popularity as a method for uncertainty quantification due to its validity guarantee. It plays a crucial role in modern decision-making processes, particularly in domains where understanding the joint distribution of random variables is essential. Under the exchangeability assumptions on the calibration and test samples, conformal prediction can provide prediction sets with valid coverage guarantees. Existing works have studied prediction intervals for single targets (Lei et al. 2018; Romano, Patterson, and Candes 2019; Sesia and Romano 2021). Given a black box algorithm to estimate the target label, current conformalized algorithms can provide prediction intervals for the target with reliable coverage probability. However, although numerous studies have developed the methodology of conformal prediction, they are limited to predicting a single label, leaving a gap in finding the prediction set for the sum of labels.
This gap is particularly relevant in applications such as transductive conformal prediction on traffic networks. For example, existing Graph Neural Network (GNN) methods can predict the label of each road, where the label can be considered as the cost of traversing that road. This problem has been studied in (Huang et al. 2024; Zargarbashi, Antonelli, and Bojchevski 2023; Zhao, Kang, and Cheng 2024). Conformalized GNN can output a prediction set of each edge’s label, but the cost of a route, which is the sum of the labels of the edges on the route, cannot be directly obtained by applying conformal prediction to individual edges. This challenge arises from two main aspects. First, the sum or average of random variables involves the convolution of the density function, and simple interval arithmetic, such as adding up the lower and upper bounds, cannot provide a confidence interval with the desired coverage. Second, the coverage of each confidence interval is in a marginal sense, so the coverage of two labels from two confidence intervals are dependent events. Consequently, it is difficult to use the conformal prediction set for a single label to devise a prediction set for multiple labels. To address this critical gap in the current literature on conformal prediction and expand its applicability to a broader range of uncertainty quantification problems, we introduce the method Conformal Interval Arithmetic (CIA), specifically designed to estimate the average or other symmetric functions of unknown labels over a certain index set, demonstrating the usefulness of our problem setting in many applications where people are interested in obtaining estimates about multiple labels.
@@ 1 Introduction
In addition to the methodology of symmetric calibration and the validity guarantee discussed above, we will introduce some important extensions of our method to make it more widely applicable and efficient. First, symmetric calibration is easy to understand when the groups of indices are disjoint. We relax this condition to allow for small amounts of overlap between groups. Second, given prior knowledge of the number of unknown labels being summed, we can stratify the prediction into different levels. Finally, we can incorporate other conformal prediction methods, such as conformalized quantile regression (Romano, Patterson, and Candes 2019), to improve prediction efficiency.
@@
  3. 3.
    We apply our methods on datasets in (Sesia and Candès 2020), and show the reliability of our method. The code of our method will be published as open source.
2 Preliminary and Problem Setup
@@
2.1 Split Conformal Prediction
@@
We will first focus on this choice and then extend it to other choices in later sections. For data in $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{cal}}}$, we evaluate $s_{i}=|y_{i}-\widehat{y}_{i}|$ and let $Q_{1-\alpha}$ be the $\lceil(1+|\mathcal{I}_{\textrm{cal}}|)(1-\alpha)\rceil$-th smallest value of $|y_{i}-\widehat{y}_{i}|$ for $i\in\mathcal{I}_{\textrm{cal}}$ [^1] If $\lceil(1+|\mathcal{I}_{\textrm{cal}}|)(1-\alpha)\rceil>|\mathcal{I}_{\textrm{cal}}|$, then $Q_{1-\alpha}=\infty$. Here, $Q_{1-\alpha}$ represents the quantile of the scores. The meaning of $Q_{1-\alpha}$ may vary depending on how the scores are defined in different sections.. Then, for $i\in\mathcal{I}_{\textrm{test}}$, the prediction set of $y_{i}$ is defined as
@@
2.2 Problem Setup and a Naive Approach
@@
$\mathbb{P}(Y_{i}+Y_{j}\in[L_{i}+L_{j},U_{i}+U_{j}])$
$\geq\mathbb{P}(Y_{i}\in[L_{i},U_{i}]\text{ and }Y_{j}\in[L_{j},U_{j}])$
$\geq 1-2\alpha.$
@@
3 Conformal Interval Arithmetic
@@
$s_{k}:=s(z_{k}):=\left|\sum_{i\in S_{k}}(y_{i}-\widehat{y}_{i})\right|,k\in[K]$
@@
$\left[\sum_{i\in S_{K+1}}\widehat{y}_{i}-Q_{1-\alpha},\sum_{i\in S_{K+1}}\widehat{y}_{i}+Q_{1-\alpha}\right].$ (1)
@@
$\mathbb{P}(Z_{1},\dots,Z_{K+1})=\mathbb{P}(Z_{\pi(1)},\dots,Z_{\pi(K+1)}),$ (2)
@@ 4 Symmetric Calibration
$S^{\textrm{cal}}_{k}:=S_{k}\cap\mathcal{I}_{\textrm{cal}}\quad\text{ and }\quad S^{\textrm{test}}_{k}=S_{k}\cap\mathcal{I}_{\textrm{test}}.$ (3)
Our task becomes finding the prediction set for $\sum_{i\in S^{\textrm{test}}_{k}}y_{i}$ for every fixed $k\in[K+1]$. To simplify the presentation, let us focus on the prediction set of $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$. Other prediction sets and the coverage guarantee in the theorems can be obtained in the same manner. Using conformal interval arithmetic, we define the score function similar to (3):
$s^{\textrm{cal}}_{k}=\left|\sum_{i\in S^{\textrm{cal}}_{k}}(y_{i}-\widehat{y}_{i})\right|$ (4)
Let $Q_{1-\alpha}$ be the $\lceil(1+K)(1-\alpha)\rceil$-th smallest value of $s^{\textrm{cal}}_{k},k=1,\dots,K$. Then we define the prediction set for $\sum_{i\in S^{\textrm{test}}_{k}}y_{i}$, denoted by $\mathcal{C}(S^{\textrm{test}}_{k})$:
$\left[\sum_{i\in S^{\textrm{test}}_{k}}\widehat{y}_{i}-Q_{1-\alpha},\sum_{i\in S^{\textrm{test}}_{k}}\widehat{y}_{i}+Q_{1-\alpha}\right].$ (5)
@@
1:  Input: labeled samples $\subscripttest\{(x_{i},y_{i})\}_{i\in\mathcal{I}\backslash\mathcal{I}_{\textrm{test}}}$, unlabeled samples $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$, a black box algorithm $\mathcal{B}$, disjoint index sets $S_{1},\dots,S_{K}\subseteq\mathcal{I}$.
2:  Output: Prediction sets for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$.
@@
4:  $\widehat{f}\leftarrow\mathcal{B}(\{(x_{i},y_{i})\}_{i\in\mathcal{I}\backslash(\mathcal{I}_{\textrm{cal}}\cup\mathcal{I}_{\textrm{test}})}$.
@@
6:  $s^{\textrm{cal}}_{k}=\left|\sum_{i\in S^{\textrm{cal}}_{k}}(y_{i}-\widehat{f}(x_{i}))\right|$ for $k\in[K]$
7:  $Q_{1-\alpha}\leftarrow\lceil(1+K)(1-\alpha)\rceil$-th smallest value of $\{s^{\textrm{cal}}_{k}:k\in[K]\}\cup\{\infty\}$.
@@
We will analyze the coverage probability of $\mathcal{C}(S^{\textrm{test}}_{K+1})$. In this section, we assume that $S_{1},\dots,S_{K},S_{K+1}$ are disjoint. We will introduce the reasoning behind this in two steps. Firstly, under the assumption of group exchangeability, the tuple of score functions $(s^{\textrm{cal}}_{1},\dots,s^{\textrm{cal}}_{K},s^{\textrm{cal}}_{K+1})$ is exchangeable. $s^{\textrm{cal}}_{k}$ has a uniform rank over $[K+1]$, so $s^{\textrm{cal}}_{K+1}\leq Q_{1-\alpha}$ with probability at least $1-\alpha$. Equivalently, this suggests that the interval
$\left[\sum_{i\in S^{\textrm{cal}}_{K+1}}\widehat{y}_{i}-Q_{1-\alpha},\sum_{i\in S^{\textrm{cal}}_{K+1}}\widehat{y}_{i}+Q_{1-\alpha}\right]$ (6)
@@
$\mathbb{P}(i\in\mathcal{I}_{\textrm{cal}})=\mathbb{P}(i\in\mathcal{I}_{\textrm{test}})=0.5$ (7)
@@
Figure 1: Diagram illustrating the partition of data into calibration and test sets for Proposition 1 and Theorem 1. The sums within each of $S_{1},\dots,S_{K},S_{K+1}$ (Top) serve as the calibration data to predict the sum within $S_{K+1}$, mirroring the calibration row for the theorem’s diagram (Bottom). As $S^{\textrm{cal}}_{K+1}$ and $S^{\textrm{test}}_{K+1}$ are exchangeable, their sums have the same distribution. Therefore, the prediction set (6) can be derived similarly to (5).
@@
The random splitting assumption (7) is slightly different from the procedure of holding out a calibration set $\mathcal{I}_{\textrm{cal}}$ with the same size as $\mathcal{I}_{\textrm{test}}$ in Algorithm 1. Theorem 1 suggests that it only requires the calibration set and test set have roughly the same size. Suppose $\mathcal{I}_{\textrm{cal}}$ and $\mathcal{I}_{\textrm{test}}$ are partitions of $\mathcal{I}_{\textrm{cal}}\cup\mathcal{I}_{\textrm{test}}$ with equal size, this complicates the proof of the theorem, as we can only swap $S^{\textrm{cal}}_{k}$ and $S^{\textrm{test}}_{k}$ when they have the same size.
5 Extensions
In this section, we will extend the applicability of the proposed method to handle overlapping subsets. Furthermore, we will introduce a stratified technique and utilize Conformal Quantile Regression (CQR) (Romano, Patterson, and Candes 2019) to enhance the efficiency of our approach.
5.1 Overlapping Subsets
Algorithm 1 can be applied to overlapping subsets $S_{1},\dots,S_{K},S_{K+1}$, but the conclusion of Theorem 1 is valid only if the subsets are disjoint. There are applications where we must assume that the subsets overlap. We will study these applications in Section 7.2. Theoretically, overlapping subsets affect the validity of the method. The following theorem shows that as long as each subset overlaps with only a small portion of the other subsets, the effect on the coverage probability is acceptable.
Theorem 2.
We still assume condition (b) and (c) in Theorem 1, and suppose that
@@ 5.1 Overlapping Subsets
Then for fixed $k\in[K+1]$, the coverage probability satisfies
@@
5.2 Stratified Conformal Prediction
Given the known number of unknown labels in each test set, we can leverage this information to enhance the efficiency of conformal prediction. For instance, consider a case where $|S^{\textrm{test}}_{k}|$ is relatively small, such as $|S^{\textrm{test}}_{k}|=1$. In this case, it is more appropriate to compare the residual with other residuals that have a smaller number of unknown labels in the calibration set. By incorporating this idea, we can modify Algorithm 1 and derive a stratified version of the algorithm that takes advantage of this additional information to improve its efficiency.
@@ 5.2 Stratified Conformal Prediction
1:  Input: labels and predictions $\subscripttrain\{(y_{i},\widehat{y}_{i})\}_{i\in\mathcal{I}\backslash\mathcal{I}_{\textrm{train}}}$, index sets $S_{1},\dots,S_{K},S_{K+1}\subseteq\mathcal{I}$. Partition of positive integers $C_{1},C_{2},\dots$
2:  Output: Prediction sets for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$.
3:  $s^{\textrm{cal}}_{k}\leftarrow\left|\sum_{i\in S^{\textrm{cal}}_{k}}(y_{i}-\widehat{y}_{i})\right|$ for $k\in[K+1]$.
4:  Find $C_{j}$ such that $|S^{\textrm{test}}_{k}|\in C_{j}$.
@@
6:  $Q^{j}_{1-\alpha}\leftarrow\lceil(1+n_{j})(1-\alpha)\rceil$-th smallest value of $\{s^{\textrm{cal}}_{k}:|S^{\textrm{cal}}_{k}|\in C_{j},k\in[K]\}\cup\{\infty\}$.
@@
5.3 Conformal Quantile Regression
From the theorems, we can observe that the form of the score function can be modified. For instance, we can adopt Conformal Quantile Regression (CQR) (Romano, Patterson, and Candes 2019). The score function can be defined as
$s^{\textrm{cal}}_{k}=\max\left\{\sum_{i\in S^{\textrm{cal}}_{k}}(\widehat{y}_{i}^{\alpha/2}-y_{i}),\sum_{i\in S^{\textrm{cal}}_{k}}(y_{i}-\widehat{y}^{1-\alpha/2}_{i})\right\}$
for $k\in[K]$, where $\widehat{y}^{\alpha/2}_{i}$ and $\widehat{y}^{1-\alpha/2}_{i}$ are the outputs of quantile regression at quantiles $\alpha/2$ and $21-\alpha/$ 2, respectively, to predict $y_{i}$. We can obtain $Q_{1-\alpha}$ in Algorithm 1 using the score functions of CQR. The prediction set $\mathcal{C}(S^{\textrm{test}}_{k})$ is then given by
$\left[\sum_{i\in S^{\textrm{test}}_{k}}\widehat{y}^{\alpha/2}_{i}-Q_{1-\alpha},\sum_{i\in S^{\textrm{test}}_{k}}\widehat{y}^{1-\alpha/2}_{i}+Q_{1-\alpha}\right].$